At 25 °C only 0.0860 mol of the generic salt AB2 is soluble in 1.00 L of water. What is the Ksp of the salt at 25 °C?

AB2(s)<------>A2+(aq)+2B-(aq)

Ksp=?
Answer #1

AB2(s)<------>A2+(aq)+2B-(aq)

Ksp = [A+2][B-]^2

[A+2] = s

[B-] = 2s

then

Ksp = (0.086)(2*0.086)^2 = 0.0025442